# SHE-Net: Syntax-Hierarchy-Enhanced Text-Video Retrieval

### v2

@@ I INTRODUCTION
Figure 2: The overview of our methods and main contributions.
@@ II-B CLIP-based Text-Video Retrieval
The remarkable success of CLIP [17] has proven effective in numerous downstream tasks, including image-text understanding [29, 30], video action recognition [31, 32], video understanding [33, 34, 35, 36, 37, 38, 39], etc. CLIP has notably contributed to the advancement of video retrieval tasks, as evidenced by recent works like CLIP4Clip [3], TS2-Net [2], STAN [7], MULTI [40] and others   [41, 42]. Videos, unlike images, contain abundant spatio-temporal information, presenting unique challenges for adaptation. While numerous methods, including CLIP4clip [3], DRL [5], X-CLIP [6] and others, attempt to enhance the CLIP architecture with temporal modelling, they typically yield only incremental improvements. Consequently, their ability to generate robust video representations remains constrained. Towards this end, CLIP-ViP [43] incorporates temporal signals into the visual encoder to better integrate the rich spatio-temporal information. Different from simply leveraging the CLS token to serve as frame features, CenterCLIP [44], TS2-Net [2], etc. introduce the remaining token features to achieve more informative and robust video representations. Unlike the aforementioned methods, in our work, we abstract the massive redundant video tokens into a visual syntax hierarchy with the guidance of text descriptions, and then exploit it to enhance the quality of generated multi-modal features.
@@ III-B Text Syntax Hierarchy Building and Initialization
Specifically, given a text description ${\boldsymbol{t}}=\{q_{1},q_{2},...,q_{N_{\rm{t}}}\}$ containing $N_{t}$ words, we apply an off-the-shelf syntax analysis toolkit to obtain the parts-of-speech tag of each word, and grammatical dependencies among words. We then build a syntax tree according to the grammatical dependencies, and parse this syntax tree into a four-level syntax hierarchical (i.e., the sentence, verbs, nouns, and adjectives) structure, which follows a global-to-local architecture. In the obtained four-layer hierarchical structure, the first layer ${\boldsymbol{H}}^{1}$ has only one special overall node ${\boldsymbol{h}}^{t1}_{1}$, representing the whole text. The second layer ${\boldsymbol{H}}^{2}=\{{\boldsymbol{h}}^{2}_{1},{\boldsymbol{h}}^{2}_{2},...,{\boldsymbol{h}}^{2}_{|{\boldsymbol{H}}^{2}|}\}$ contains all verbs in sentences, and all verb nodes are connected to the overall node. Subsequently, the third layer ${\boldsymbol{H}}^{3}=\{{\boldsymbol{h}}^{3}_{1},{\boldsymbol{h}}^{3}_{2},...,{\boldsymbol{h}}^{3}_{|{\boldsymbol{H}}^{3}|}\}$ contains all entity nouns, which describe the entities involved in actions, and each node in ${\boldsymbol{H}}^{3}$ is connected to its associated verb nodes in ${\boldsymbol{H}}^{2}$. Specially, if verbs are missing from the text description, the model will fill in a special token [[[[EXIST]]]] to represent the meaning of existence, and connect it to nouns to ensure that the model can work normally. Finally, the bottom fourth layer ${\boldsymbol{H}}^{4}=\{{\boldsymbol{h}}^{4}_{1},{\boldsymbol{h}}^{4}_{2},...,{\boldsymbol{h}}^{4}_{|{\boldsymbol{H}}^{4}|}\}$ is the adjectives, connecting to their related entity noun nodes.
@@ III-C Video Syntax Hierarchy Building and Feature Fusion
Videos are natural signals with significant spatial and temporal redundancy. In existing literature [11, 9], the methods for obtaining video features have typically utilized separate multilayer perceptron (MLP) encoders to extract features at the video, action, and entity levels from frame-level features, which have limited capability in aggregating informative video features. In this paper, we propose leveraging the informative and hierarchical structure of text to guide the fusion of video features at varying granularities. Considering that videos exhibit a hierarchical structure analogous to text, we aim to enhance video representations and filter out redundant signals across granularity levels. To this end, we decompose videos into three hierarchical components, namely video content, frames (i.e., actions), and patches (i.e., entities), establishing a three-tier video syntax hierarchy in the sequence of ”video content-actions-entities”. Through this approach, we employ distinct analytical methods tailored to each granularity level to enhance the extracted features.
@@ III-D Syntax-Hierarchy-Enhanced Similarity Calculation
After obtaining the visual hierarchy corresponding to the text hierarchy, we design a syntax-hierarchy-enhanced similarity computation to obtain the final text-video similarity. Unlike previous work, which usually calculates the similarity between global-pooling multi-modal features or roughly calculates the similarity between the corresponding layers, we first align the established text and video syntax hierarchies and calculate their similarity scores node-by-node as follows:
@@
Then, logically, different actions have different importance to the overall text, and different entities also have different importance to the same action. Therefore, we assign text-video node pairs with different weights ${\boldsymbol{w}}^{\theta}_{i}$ based on the syntax hierarchy. Specifically, for the whole node $h^{1}_{1}$ in $H^{1}$, we directly set its weight ${\boldsymbol{w}}^{1}_{i}$ to 1. Besides, for action-level nodes, we utilize the similarity between the overall node ${\boldsymbol{h}}_{1}^{1}$ and each verb node in ${\boldsymbol{H}}^{2}$ to assign weights $w^{2}_{i}$ to each action node as follows:
@@ IV-D Comparison with State-of-the-Arts
IV-E Ablation Studies
@@
IV-E1 Effectiveness of Syntax-Hierarchy-Enhanced Modules
@@
IV-E2 The Impact of Weights in Video Syntax-Hierarchy Building
@@
IV-F Qualitative Results
@@
Additionally, to clarify further the rationale behind the construction of video syntax structure, we present some examples of matching visual features for verbs in Fig.4. These examples illustrate the alignment of verbs with frame sequences, highlighting that different verbs are associated with specific frames within the sequence. Our method retrieves the most $\lambda_{\rm{frame}}$-relevant frames for each verb to construct the visual action features that correspond to the verb, thus mitigating the impact of noise.

### v3

@@ I INTRODUCTION
Figure 2: The overview of the proposed SHE-Net method and main contributions. (a) SHE-Net first constructs and utilizes the text syntax hierarchy to learn the grammatical structure of the given text. (b) This method then develops the video syntax hierarchy, guided by the corresponding text syntax, effectively aggregating numerous video signals into information-dense. (c) Finally, by leveraging these two syntax hierarchies jointly, this method enhances the cross-modal alignment across various levels of granularity, thus achieving superior retrieval performance.
@@ II-B CLIP-based Text-Video Retrieval
The remarkable success of CLIP [17] has proven effective in numerous downstream tasks, including image-text understanding [29, 30, 31], video action recognition [32, 33], video understanding [34, 35, 36, 37, 38, 39, 40], etc. CLIP has notably contributed to the advancement of video retrieval tasks, as evidenced by recent works like CLIP4Clip [3], TS2-Net [2], STAN [7], MULTI [41] and others   [42, 43, 44, 45]. Videos, unlike images, contain abundant spatio-temporal information, presenting unique challenges for adaptation. While numerous methods, including CLIP4clip [3], DRL [5], X-CLIP [6] and others, attempt to enhance the CLIP architecture with temporal modelling, they typically yield only incremental improvements. Consequently, their ability to generate robust video representations remains constrained. Towards this end, CLIP-ViP [46] incorporates temporal signals into the visual encoder to better integrate the rich spatio-temporal information. Different from simply leveraging the CLS token to serve as frame features, CenterCLIP [47], TS2-Net [2], etc. introduce the remaining token features to achieve more informative and robust video representations. Unlike the aforementioned methods, in our work, we abstract the massive redundant video tokens into a visual syntax hierarchy with the guidance of text descriptions, and then exploit it to enhance the quality of generated multi-modal features.
@@ III-B Text Syntax Hierarchy Building and Initialization
Specifically, given a text description ${\boldsymbol{t}}=\{q_{1},q_{2},...,q_{N_{\rm{t}}}\}$ containing $N_{t}$ words, we apply an off-the-shelf syntax analysis toolkit to obtain the parts-of-speech tag of each word, and grammatical dependencies among words. We then build a syntax tree according to the grammatical dependencies, and parse this syntax tree into a four-level syntax hierarchical (i.e., the sentence, verbs, nouns, and adjectives) structure, which follows a global-to-local architecture. In the obtained four-layer hierarchical structure, the first layer ${\boldsymbol{H}}^{1}$ has only one special overall node ${\boldsymbol{h}}^{t1}_{1}$, representing the whole text. The second layer ${\boldsymbol{H}}^{2}=\{{\boldsymbol{h}}^{2}_{1},{\boldsymbol{h}}^{2}_{2},...,{\boldsymbol{h}}^{2}_{|{\boldsymbol{H}}^{2}|}\}$ contains all verbs in sentences, and all verb nodes are connected to the overall node. Subsequently, the third layer ${\boldsymbol{H}}^{3}=\{{\boldsymbol{h}}^{3}_{1},{\boldsymbol{h}}^{3}_{2},...,{\boldsymbol{h}}^{3}_{|{\boldsymbol{H}}^{3}|}\}$ contains all entity nouns, which describe the entities involved in actions, and each node in ${\boldsymbol{H}}^{3}$ is connected to its associated verb nodes in ${\boldsymbol{H}}^{2}$. Specially, in instances where verbs are absent or when the analysis tools cannot establish these relationships, we introduce a special token, [[[[EXIST]]]]. This token symbolizes the notion of existence and links it to all words in the sentence. This mechanism helps maintain the model’s functional efficacy by preserving semantic coherence, even in incomplete syntactic structures. Finally, the bottom fourth layer ${\boldsymbol{H}}^{4}=\{{\boldsymbol{h}}^{4}_{1},{\boldsymbol{h}}^{4}_{2},...,{\boldsymbol{h}}^{4}_{|{\boldsymbol{H}}^{4}|}\}$ is the adjectives, connecting to their related entity noun nodes.
@@ III-C Video Syntax Hierarchy Building and Feature Fusion
Videos are natural signals with significant spatial and temporal redundancy. In existing literature [11, 9], the methods for obtaining video features have typically utilized separate multilayer perceptron (MLP) encoders to extract features at the video, action, and entity levels from frame-level features, which have limited capability in aggregating informative video features. In this paper, we propose leveraging the informative and hierarchical structure of text to guide the fusion of video features at varying granularities. Considering that videos exhibit a hierarchical structure analogous to text, we aim to enhance video representations and filter out redundant signals across granularity levels. To this end, we decompose videos into three hierarchical components, namely video content, frames (i.e., actions), and patches (i.e., entities), establishing a three-tier video syntax hierarchy in the sequence of “video content-actions-entities”. Through this approach, we employ distinct analytical methods tailored to each granularity level to enhance the extracted features.
@@ III-D Syntax-Hierarchy-Enhanced Similarity Calculation
After obtaining the visual hierarchy corresponding to the text hierarchy, we design a syntax-hierarchy-enhanced similarity computation to obtain the final text-video similarity. Unlike previous work, which usually calculates the similarity between global-pooling multi-modal features or roughly calculates the similarity between the corresponding layers, we first align the established text and video syntax hierarchical nodes and calculate their similarity scores node-by-node as follows:
@@
Then, to calculate the similarity for each layer, we employ a learnable weighted sum approach. Logically, various actions possess different importance within the overall text, and disparate entities also carry distinct significance for the same action. Therefore, we assign text-video node pairs with different learnable weights ${\boldsymbol{w}}^{\theta}_{i}$ based on the syntax hierarchy. Specifically, for the whole node $h^{1}_{1}$ in $H^{1}$, we directly set its weight ${\boldsymbol{w}}^{1}_{i}$ to 1. Besides, for action-level nodes, we utilize the similarity between the overall node ${\boldsymbol{h}}_{1}^{1}$ and each verb node in ${\boldsymbol{H}}^{2}$ to assign weights $w^{2}_{i}$ to each action node as follows:
@@ IV-D Comparison with State-of-the-Arts
IV-E Training and Inference Time Consumption
During the training phase, our model results in an increase in training time by approximately 15% compared to the baseline model, TS2Net. The computational processes of the model primarily involve extracting textual and visual features, constructing textual syntactic structures, fusing temporal features derived from frame features, and executing syntax-enhanced visual feature fusion and similarity calculations. The first three components can be pre-computed offline and stored for inference, thus significantly reducing computational time during inference. Consequently, our proposed modules incur an additional 5% in computational cost during inference relative to that of the baseline model, TS2Net. Our method demonstrates a significant improvement in performance with a modest increase in computational time.
TABLE IX: Comparison of Text-to-Video Retrieval Performance Utilizing Visual Encoders with Varying Parameter Quantities on MSR-VTT.
[…]
ViT-L-32 48.5 77.3 86.5 2.0 11.0 212.3
ViT-L-16 54.5 79.5 88.2 1.0 9.3 222.2
@@ IV-E Training and Inference Time Consumption
IV-F Ablation Studies
IV-F1 Effectiveness of Syntax-Hierarchy-Enhanced Modules
@@
IV-F2 The Impact of Weights in Video Syntax-Hierarchy Building
@@
IV-F3 The Impact of Vision Encoders with Different Parameter Sizes
[…]
IV-G Qualitative Results
@@
Additionally, to clarify further the rationale behind the construction of video syntax structure, we present some examples of matching visual features for verbs in Fig.4. These examples illustrate the alignment of verbs with frame sequences, highlighting that different verbs are associated with distinct frames within the sequence. For instance, in the sentence “A person is singing a song, and the ladies are dancing”, the weights assigned by our methodology to “sing” and “dance” are 0.4970 and 0.5030, respectively, suggesting that these two actions hold remarkably similar importance within this context. Meanwhile, the weights for “person”, “song”, and “ladies” are 0.2970, 0.3543, and 0.3487, respectively, indicating that “song” is more prominent than “person”, while “ladies” also retain significant importance in this context. In this paper, our method retrieves the most $\lambda_{\rm{frame}}$-relevant frames for each verb to construct the visual action features that correspond to the verb, thus mitigating the impact of noise.